OpenStudy (laddiusmaximus):

tanx(1-sin^2x)=1/2sin2x

OpenStudy (anonymous):

\[1-\sin^2(x)=\cos^2(x)\] Plug in, multiply by sin(x)/cos(x) (ends up being sin(x)cos(x)) and simplify.

OpenStudy (laddiusmaximus):

I got that part. and tanx becomes sinx/cosx, which leaves me with sinx/cosx(cos^2x) whats after that? algebra?

OpenStudy (laddiusmaximus):

i end up with sinx/cosx but how does that equal 1/2sin2x?

OpenStudy (anonymous):

Is it (1/2)sin2x or 1/(2sin(2x)?

OpenStudy (laddiusmaximus):

(1/2)sin2x

OpenStudy (anonymous):

Okay, sin(x)/sin(2x) simplifies to 1/(2cos(x)).\[\sin(x)\cos(x) = (1/2)\sin(2x) \rightarrow\] Seems like if we go down this road, we're just going to end up with 1/2 = 1/2...

OpenStudy (laddiusmaximus):

whatt identity would that be?

OpenStudy (anonymous):

sin(2x) = 2sin(x)cos(x)